How many water contain in a 7 inch pipe in 1 meter depth?

Grade:12th pass

1 Answers

Adithya Venkateswaran
7 years ago
12 in = 30 cm
7 in = 30×7÷12 cm
D = 2.5×0.07 m
Therefore the volume of water is
πD²÷4×1m = 22÷7×2.5×2.5×0.07×0.07÷4 = 0.0240625m³ or 2.40625 × 10⁴ cm³